OpenStudy (anonymous):

How do I write this point (0, 2) and a line that's going horizontal? Please help

OpenStudy (anonymous):

do you need to graph it? if so, since the x coordinate is 0 and the y coordinate is 2, you would stay at 0 and just go up to spots and draw a horizontal line at 2

OpenStudy (anonymous):

No I just need it in an equation

OpenStudy (anonymous):

y=2

OpenStudy (anonymous):

Y= mx + b form

OpenStudy (tester97):

like y=mx+b kind of equation?

OpenStudy (anonymous):

Yea

OpenStudy (anonymous):

since there is no slope you would forget the mx part which would make it y=2 :)

OpenStudy (tester97):

yea i was gunna say that ;)

OpenStudy (anonymous):

great minds think alike ;)

OpenStudy (anonymous):

Its not making any sense to me tho " great minds"

OpenStudy (anonymous):

Lol

OpenStudy (anonymous):

hahaha sorry, so you needed the points to be in slope-intercept form which is y=mx+b. The points given were (0,2) which means that there is no slope, just a y intercept of 2. with that being said, points with no slopes are always horizontal, so you just draw a horizontal line at (0,2).

OpenStudy (tester97):

well for someone who this is not making sense to you sure do get it right

OpenStudy (anonymous):

No need the attitude...im in need of help so get it together :)

OpenStudy (anonymous):

X in coordinates like (x,y) is always the slope. but in your case since the X was a 0, there is absolutely no slope at all. if it would have been (3,2) then your slope would have been 3/1. get it?

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Need it as y= mx + b please

OpenStudy (anonymous):

that is the formula! y=2 is the answer lol

OpenStudy (anonymous):

K thanx!!!!

OpenStudy (tester97):

@hubbabubba i was not giving any attitude :P lol i was kinda giving a compliment

OpenStudy (anonymous):

read what i commented or whatever earlier lol. a slope of zero is no slope at all

OpenStudy (anonymous):

you're welcome, i hope it kinda made sense :)

OpenStudy (anonymous):

Alright pplz lol it did some how